DIRC Operations Weekly Report
Period: Wednesday 2007/03/01 - Wednesday 2007/03/06
Oncall: Alejandro Perez
Fast Monitoring: Alejandro Perez, Bradley Wogsland
File location: $BFROOT/www/Detector/DIRC/WeeklyReports/report_20070306.html
Plot directory: $BFROOT/www/Detector/DIRC/WeeklyReports/Plots/report_20070221
PEP peak lumi: ~8.3 x 1033
(previous week: ~8.2 x 1033)
BaBar integrated lumi: ~3.1
(previous week: ~3.2 fb-1)
Run 6 PEP delivery:~3.3
(previous week: ~3.3 fb-1)
Run 6 total delivery: 13.2 fb-1
Run 6 total BaBar recorded: 12.7 fb-1
including 0 fb-1 off-peak data
PEP activities this week:
MD on March 5. PEP was testing relationship between Background and beam parameters.
Foreseen work opportunities (PEP down periods):
3 days ROD begining on March 26. Backward door are going to be close, no expected DIRC work.
BaBar activities/issues this week:
- DIRC commissioner and ODF expert paged at 5pm on March 02 because of 80% dead time from DIRC not due to Backgrounds. Damage indicates that one (or more) DFBs in sector 11 have no conection with ROM (no indication which DFB may the problematic one because the log file get lost). With the start of new run the problem went away
logbook for details
- On March 06 at the afternoon high backgrounds produciong asymetric "shoulders" in DIRC timing plots. Emmanuele made a test with MCC, revelating that high backgrounds are due to bad trickle injection in LER.
Weekly Bkg Plot
- Test of DCC and DFBs that arrived from France last week.
- Maybe replace two DFBs in sector 1 with a group of 8 innefficient or flaky channels (require bridge).
- A few noisy channels need to be masked.
- Unmask all masked channels to find which of these are actually noisy. After December DFB work/movement masked channels may no longer be the noisy ones.
Bradley, Jose and Alejandro rotating backup & on-call commissioning duties